OpenStudy (anonymous):

A bowl contains 8 blue marbles and 5 green marbles. Elizabeth randomly draws 4 marbles from the bowl. She does not replace the marbles after each draw. What is the probability that she draws 1 blue marble and 3 green marbles? A. B. C. D.

OpenStudy (anonymous):

a.13/143 b.4/13 c.4/143 d.16/143

OpenStudy (kropot72):

\[\large P(1\ blue,\ 3\ green)=\frac{C(8, 1)\ \times C(5,\ 3)}{C(13,\ 4)}=\frac{8\times10}{715}\]

OpenStudy (anonymous):

i dont understand is that the answer? or how do you get the answer ?

OpenStudy (kropot72):

No, it is part way to the answer. You need to simplify as follows:  Can you finish it now?